prices, demand, and use of Austin Enterprises makes and sels three types of dress shirts. Management is trying to determine the most proftable mix Sales follow Basic Classic $ 38 18,000 $ 200 28,000 S 70 Sales price Maximum annual demand (units) Input requirement per unit 1,000 Direct material Direct labor 0.7 yards 05 yards 0.8 yards 8 hours 09 hours 2 hours Costs Variable costs Materia s Direct labor Factory overhead $ 18 per yard $ 14 per hour $ 5 per direct labor-hour ,0 % of sales price Annual fixed costs Marketing Administration $50,000 s 7,500 $44,000 The company faces two limits:(1) the volume of each type of shirt that it can sell see maximum annual demand) and (2) 35,500 direct labor-hours per year caused by the plant layout Required: -1. Assuming the company can satisty the annual demand, calculate the contribution margin for each type of dress shirt using the table below Total revenue

media%2F566%2F5662ff63-c4e1-44eb-b5dd-6b

media%2Fba8%2Fba8330b2-a267-4934-8312-71Please answer step E

media%2F9db%2F9dbef4ac-db08-4d7e-80f9-61
Answer #1

As per policy, only four parts of a question is allowed to answer, so answering a to d here :

1)
a1) Basic Classic Formal Total
Total revenue 684000 770000 5600000 7054000
Total variable costs 603000 594000 5219200 6416200
Contribution margin 81000 176000 380800 637800
a2) Operating profit = contribution - fixed costs
637800 - (50000+7500+44000) = 637800-101500 = 536300
b1) contribution margin for eash shirt per constrained resource ie. Direct labor:
Basic Classic Formal
T. Contribution margin 81000 176000 380800
total labour hours 18000*0.9=16200 11000*2=22000 28000*8=224000
contri margin per LH 5 8 1.7
b2) Most profitable use of constrained resource : Classic
c) Total constrained resource = 35500 LH
Less: classic = 22000
Remainder = 13500
Second Shirt of mix = Basic
So, Product mix for maximum profit = Classic and Basic
d) Contribution margin of the product mix :
Basic Classic
Annual demand 13500/0.9=15000 11000
Total revenue 15000*38=570000 11000*70=770000
Less: Variable cost:
Material 15000*0.7*18=189000 11000*0.5*18=99000
D. Labor 13500*14=189000 22000*14=308000
F. Overhead 13500*5=67500 22000*5=110000
Marketing (15000*38*10%)= 57000 (11000*70*10%)= 77000
Total variable costs 502500 594000
Contribution 67500 176000

d2) Operating profit = total contri - fixed cost = 243500 - 101500 = $142000

===============